# Configurable photonic simulator for quantum field dynamics

### v1

@@ D.2 Entropy density of the GGE and mode populations
where $n_{k}=\braket{0}{\bm{n}_{k}}{0}$ specifies the quasi-particle content of the initial state (that is conserved during time evolution).
@@
describing quasi-particle excitations. Note here that the appearance of the relativistic dispersion $\omega_{k}$ guarantees Lorentz-invariance of the momentum integral measure in the continuum.
The ground state |$\ket{\Omega}$ of (30), or, equivalently, (76), corresponds to the quasi-particle vacuum $\tilde{\bm{a}}_{k}\ket{\Omega}=0$, while the input-state vacuum $\bm{a}_{j}\ket{0}=0$ contains $n_{k}$ quasi-particles. The latter follows after expanding the quasi-particle mode operators $(\tilde{\bm{a}}_{k},\tilde{\bm{a}}_{k}^{\dagger})$ in terms of the position-space, i.e., photonic, mode operators $(\bm{a}_{j},\bm{a}_{j}^{\dagger})$, for which we combine Eqs. (77) and (70) with $\bm{a}_{j}=(\bm{\phi}_{j}+\bm{\pi}_{j})/\sqrt{2}$, leading to
@@
$\eta_{k}=\dfrac{1}{\sqrt{\epsilon\omega_{k}}}+\sqrt{\epsilon\omega_{k}},\quad\kappa_{k}=\dfrac{1}{\sqrt{\epsilon\omega_{k}}}-\sqrt{\epsilon\omega_{k}}.$ (79)
@@
$\begin{split}n_{k}&=\braket{0}{\tilde{\bm{a}}_{k}^{\dagger}\tilde{\bm{a}}_{k}}{0}\\ &=\dfrac{\kappa_{k}^{2}}{4N}\sum_{j,j^{\prime}\in\mathcal{J}}e^{-2\pi i(j-j^{\prime})k/N}\langle 0|\bm{a}_{j^{\prime}}\bm{a}_{j}^{\dagger}|0\rangle\\ &=\dfrac{\kappa_{k}^{2}}{4}\\ &=\dfrac{1}{4}\left(\dfrac{1}{\epsilon\omega_{k}}+\epsilon\omega_{k}-2\right).\end{split}$ (80)

### v2

@@ D.2 Entropy density of the GGE and mode populations
where $n_{k}=\braket{0|\bm{n}_{k}|0}$ specifies the quasi-particle content of the initial state (that is conserved during time evolution).
@@
describing quasi-particle excitations. Note here that the appearance of the relativistic dispersion $\omega_{k}$ guarantees Lorentz-invariance of the momentum integral measure in the continuum and necessitates squeezing.
[…]
@@
$\eta_{k}=\sqrt{\epsilon\omega_{k}}+\dfrac{1}{\sqrt{\epsilon\omega_{k}}},\quad\kappa_{k}=\sqrt{\epsilon\omega_{k}}-\dfrac{1}{\sqrt{\epsilon\omega_{k}}}.$ (100)
@@
$\begin{split}n_{k}&=\braket{0|\tilde{\bm{a}}_{k}^{\dagger}\tilde{\bm{a}}_{k}|0}\\ &=\dfrac{\kappa_{k}^{2}}{4N}\sum_{j,j^{\prime}\in\mathcal{J}}e^{-2\pi i(j-j^{\prime})k/N}\langle 0|\bm{a}_{j^{\prime}}\bm{a}_{j}^{\dagger}|0\rangle\\ &=\dfrac{\kappa_{k}^{2}}{4}\\ &=\dfrac{1}{4}\left(\epsilon\omega_{k}+\dfrac{1}{\epsilon\omega_{k}}-2\right).\end{split}$ (101)